OpenStudy (anonymous):

Leah purchased a lawn mower for $3,865. It depreciates 1.5% each year. What is the value of the lawn mower after seven years? $3,807.03 $4,289.55 $3,831.31 $3,476.99

OpenStudy (johnweldon1993):

1 more :P same equation as that last one Vfinal = 3865(1-r)^t Vfinal = 3865(1-.015)^7 =3476.987 okay now i say my goodbyes :)

OpenStudy (anonymous):

youre so nice, thanks:)

OpenStudy (johnweldon1993):

anytime :)